Fermi National Laboratory
 

Accelerator Update for June 22, 2007 - June 25, 2007

 

Friday June 22

The day shift began with Operations monitoring store 5506 (58.86E30), the Recycler maintaining a stash (277.0E10), with the Antiproton Source (Pbar) stacking (32.3E10), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI and MiniBooNE taking beam.

At 12:35 PM, the NuMI horn tripped off on a ground fault indication. Operators followed the procedure to burn off the flakes. NuMI beam resumed at 1:31 PM.

Operations reported at 3:58 PM that the Recycler had stashed antiprotons two times during the day shift.

Operations terminated store 5506 at 4:17 PM.

The Tevatron (TeV) went into shot setup at 4:20 PM.

The Tev quenched at 6:05 PM.

D0 is making a short 2 hour access beginning at 8:18 PM.

D0 access is complete at 10:08 PM.

Operations reported at 11:56 PM that the Recycler had stashed antiprotons three times during the evening shift.

Saturday June 23

The midnight shift began with studies in the TeV, and the stash at 185.9E10, with Pbar stacking (16.7E10), and with NuMI and MiniBooNE taking beam.

The Tevatron (TeV) went into shot setup at 3:24 AM.

Operations established store 5510 at 6:09 AM with an initial luminosity of 140.08 E30.

Operations reported at 8:06 AM that the Recycler had stashed antiprotons two times during the midnight shift.

Operations reported at 7:58 PM that the Recycler had stashed antiprotons five times during the day shift.

Sunday June 24

The midnight shift began with Operations monitoring store 5510 and the stash , with Pbar stacking, and with NuMI and MiniBooNE taking beam.

Operations terminated store 5510 at 12:05 AM.

At 12:13 AM, Main Injector RF station 17 (MIRF17) lost its bias power. The MI is now running without MIRF17 which will be looked at on Monday.

The Tevatron (TeV) went into shot setup at 12:43 AM.

Operations established store 5511 at 3:32 AM with an initial luminosity of 192.11 E30.

Operations reported at 8:06 AM that the Recycler had stashed antiprotons four times during the midnight shift.

Operations reported at 7:59 PM that the Recycler had stashed antiprotons five times during the day shift.

Monday June 25

The midnight shift began with Operations monitoring store 5511 and the stash , with Pbar stacking, and with NuMI and MiniBooNE taking beam.

A vacuum burst occurred in the AP2 line at 12:28 AM.

At 1:00 AM an access was made into the PBar Transport Enclosure to look for a water leak and repair the AP2 vacuum. No beam to PBar, NuMI and MiniBooNE while the access is being made.

The AP2 water leak was fixed by 2:30 AM. Vacuum leak repair is still in progess.

Operations terminated store 5511 at 6:57 AM.

D0 and CDF went into controlled access at 7:12 AM, D0 will be in access for 10 hours to repair some of their calorimeter power supplies.

The Plans for Monday

Repair the AP2 vacuum, use this downtime to make repairs in other machines, put a store in the TeV when the D0 access is completed.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/25/2007    email Fermilab